Month Jan. Feb. Mar. Apr.
Actual sales 100 80 90 95
Forecasted sales 95 70 95 80

Inputs will be exact numbers.

  1. Use the above data to determine the MAD value.
  1. Assume the MAD value from the table above is 1, use the assumed MAD value 1, please calculate the TS value.
Answer #1

Mean absolute deviation (MAD) is the average of the absolute differences between actual and forecasted values.

Hence, MAD = (|100 - 95| + |80 - 70| + |90 - 95| + |95 - 80|) / 4

=> MAD = (5 + 10 + 5 + 15) / 4

=> MAD = 35/4 = 8.75

Tracking signal (TS) is the ratio of cumulative error and MAD.

Cumulative error = (100 - 95) + (80 - 70) + (90 - 95) + ( 95 - 80) = 5 + 10 + (-5) + 15 = 25

Assumed MAD value = 1

=> TS = Cumulative error / assumed MAD = 25/1 = 25